OpenStudy (anonymous):

could anyone help me with this question, please? How can I get eigenfunctions and eigenvalues for a particle in a one-dimensional box of a length a from time-independent Schrodinger equation?

OpenStudy (michele_laino):

hint: we have to solve thie ODE: \[\huge \frac{{ - {\hbar ^2}}}{{2m}}\frac{{{d^2}\psi }}{{d{x^2}}} = E\psi \] with these boundary conditions: \[\huge \psi \left( 0 \right) = \psi \left( a \right) = 0\]

OpenStudy (michele_laino):

oops.. this* ODE...

OpenStudy (anonymous):

hello Michele_Laino, I am trying to solve it, if i have some problems , i will let you know, but what do you mean bu ODE

OpenStudy (anonymous):

I did it, and I found the answer Thank you so much

OpenStudy (michele_laino):

ODE, stands for \(O\)rdinary \(D\)ifferential \(E\)quation